OpenStudy (aayushi.somani):

integrate by parts : x tan ^-1 x @SolomonZelman

OpenStudy (aayushi.somani):

i ended up here : x^2 /2 tan^-1 x - § 1 / root (x^2 + 1) * x^2 /2 .dx and i lost it..

OpenStudy (solomonzelman):

this is very nasty for me to explain, because I am myself a student... § x tan \(\scriptsize\color{black}{ ^{-1}}\)x dx I don't see any other method to use other than your's :) tan \(\scriptsize\color{black}{ ^{-1}}\)x=f df = 1/(x²+1) dx dg=xdx g=x²/2 and this all is in `§f dg= fg - §g df `

OpenStudy (solomonzelman):

This is how I would start at least.

OpenStudy (aayushi.somani):

m sorry but here we need to solve it "by parts" and not by substitution!

OpenStudy (solomonzelman):

No, I am integrating ' x tan \(\scriptsize\color{black}{ ^{-1}}\)x ' like this, and it is by parts.

OpenStudy (aayushi.somani):

§f1 . f2 .dx= f1 §f2 dx - § [df1/dx §f2 dx ] dx i meant solve it by this!

OpenStudy (solomonzelman):

I am very confusing often times, but when I am continuing I am getting, ½ x² tan \(\scriptsize\color{black}{ ^{-1}}\)x - § x² / [ 2(x² +1) ] dx ½ x² tan \(\scriptsize\color{black}{ ^{-1}}\)x - ½§ x² / [ (x² +1) ] dx then, long division for the second integral, ½ x² tan \(\scriptsize\color{black}{ ^{-1}}\)x - ½§ 1- [1/(x² +1) ] dx

OpenStudy (solomonzelman):

½ x² tan \(\scriptsize\color{black}{ ^{-1}}\)x - ½§ 1/(x² +1) dx - ½ § 1 dx

OpenStudy (solomonzelman):

½ x² tan\(\scriptsize\color{black}{ ^{-1}}\)(x) + ½ tan\(\scriptsize\color{black}{ ^{-1}}\)(x) -x/2 +C

OpenStudy (solomonzelman):

then factor out of 1/2

OpenStudy (aayushi.somani):

is the differentiation of tan^-1 x = x^2 +1 or root ( x^2 +1 ) ?

OpenStudy (solomonzelman):

none

OpenStudy (solomonzelman):

1/ (x² +1)

OpenStudy (solomonzelman):

yeah, .. I am not a very good helper.... igtg

OpenStudy (aayushi.somani):

ohk.. thankyou soo much this was my mistake..

OpenStudy (aayushi.somani):

okay thanks.. i got the answer (: and ur a great great helper :D

OpenStudy (solomonzelman):

yw... btw, the highest math course I took was trigonometry, will leave it with this... bye and yw.

OpenStudy (aayushi.somani):

bbye (: @SolomonZelman